# Semiparametric KSD test: unifying score and distance-based approaches for goodness-of-fit testing

### v1

@@ Appendix E Preliminaries on stable convergence
Classical weak convergence results are established for the convergence of the quantile function, which is not sufficient for our purpose in this conditional convergence scenario, and a stronger notion of weak convergence is needed. To this end, we apply the notion of stable convergence [hausler2015stable], which is a stronger form of weak convergence and is suitable for studying the asymptotic behavior of conditional distributions.
@@
Generally speaking, stable convergence requires the weak limit of the sequence $Y_{n}$ keeps “stable” when conditioning on different sub-$\sigma$-algebra $\mathcal{G}$. The following lemma, adapted from Theorem 3.2 in [hausler2015stable], provides an equivalent characterization of stable convergence from this perspective.
@@
A key technical result applied in the proofs is the stable version of the central limit theorem. The following lemma, adapted from Theorem 6.1 in [hausler2015stable], states the conditions under which a martingale difference array converges stably to a normal distribution:
@@ F.1 Proof of Theorem 1
Under the regularity conditions, the law of large numbers for U-statistics [van2000asymptotic, Theorem 12.3] implies that
@@ F.2 Proof of Theorem 2
Theorem 2 can be proved by following the strategy mentioned in the main text. We refer the audience to the proofs of Theorem 3 or Corollary 2 as natural extensions of this proof strategy.
@@ F.3 Proof of Theorem 3
Our proof relies on the concept of regular conditional distribution and convergence of Markov kernel, especially stable convergence. We provide introduction and basic results on the concept of stable convergence in Appendix E. See [hausler2015stable] for a detailed introduction to this framework.
We prove the desired stable convergence result through the following lemma, which is a generalization of [fernandez2024general, Theorem 1 ]. We first introduce some notation. Define $\mathcal{F}=\sigma(\cup_{n=1}^{\infty}\mathcal{F}_{n})$ to be the smallest $\sigma$-algebra generated by the whole data sequence. Notice that $\widetilde{T}_{n}$ and $\widetilde{S}_{n}$ both depend on the data through $\hat{\theta}_{n}$, which only depends $\{X_{i}\}_{i=1}^{n}$. Based on this observation, deriving the convergence results conditioning on $\mathcal{F}$ is equivalent to conditioning on $\mathcal{F}_{n}$. Thus, for simplicity, we work on the $\mathcal{F}$-stable convergence of the test statistic in the proof in this part.
@@
We use the definition of stable convergence in [hausler2015stable, Page 6 ] to prove it. Recall Definition of the operator norm in Eqn. (30). Let us choose
@@ F.4 Proof of Theorem 4
We use the results in Theorem 2 and 3 combining with [niu2022reconciling, Lemma 3 ] to prove the asymptotic validity of the test. We state their lemma here for completeness.
@@
which is equivalent to $n\widetilde{T}_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$ with notation used in [niu2022reconciling]. Applying Lemma 19, we have
@@
where in the first inequality, we use the fact that $\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L})\geq\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})=2\varepsilon$. The last expression can be bounded by the following lemma, which is indeed a stronger result. The proof of this result can be found in [matsubara2022robust, Lemma 2 ].
@@ F.5 Proof of Proposition 2
We will verify the conditions in the Kolmogorov’s three series theorem [durrett2019probability, Page 150]. By choosing the truncation threshold $A$ in the theorem to be $\infty$, the first condition in the three series holds automatically. Thus, it suffices to show that
@@ F.6 Proof of Corollary 2
We directly prove Eqn. (37) in this part. [fernandez2024general, Theorem 1 ] is used to prove this result. We state it here for convenience.
@@
still hold for any $\varepsilon>0$. Then by standard Lindeberg-Feller Central Limit Theorem [durrett2019probability, Theorem 3.4.10 ], we verify Condition (C2.1).
@@ F.7 Verification of the conditions in Lemma 5
We will mainly use a slightly modified version of [fernandez2024general, Proposition 6 ]. The version of usage is stated as below:
@@ G.2 Verification of conditions in Lemma 14
By Le Cam’s third lemma [van2000asymptotic, Theorem 6.6], we conclude that under local alternative $\mathcal{L}_{n}$,
@@ G.3 Proof of the additive case in Theorem 5
$\mu_{i}$ $=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})],$
@@
$p_{1,n}(x)$ $=\left(1-\frac{\gamma}{\sqrt{n}}\right)p_{\theta_{0}}(x)+\frac{\gamma}{\sqrt{n}}g(x)$
$=p_{\theta_{0}}(x)\left(1+\frac{\gamma}{\sqrt{n}}\left(\frac{g(x)}{p_{\theta_{0}}(x)}-1\right)\right),$
which can be fitted into the multiplicative local alternatives with $h(x)=\gamma(g(x)/p_{\theta_{0}}(x)-1)$. It can be directly checked that Assumption 2 hold. Thus, by Theorem 5, we immediately have Theorem 5 holds with
$\mu_{i}$ $=\mathbb{E}_{\theta_{0}}\left[S^{*}(\phi_{i},X,\theta_{0})\cdot\gamma\left(\frac{g(x)}{p_{\theta_{0}}(x)}-1\right)\right]$
$=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})]-\gamma\mathbb{E}_{\theta_{0}}[S^{*}(\phi_{i},X,\theta_{0})]$
$=\gamma\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})],$
@@ H.1 Proof of Lemma 3
Notice that $X\sim\mathbb{P}_{\theta}$, the property of kernelized Stein class in Eqn. (6), we have
@@ H.2 Proof of Lemma 4
We begin by explicitly calculate the left hand side of Equation (73). It is straightforward to check that $\widetilde{S}_{n}(f)$ is a linear functional on $\mathcal{H}^{d}$ By Riesz representation theorem [yosida2012functional, Page 81 ], there exists a unique $f_{n}^{\star}\in\mathcal{H}^{d}$ such that $\widetilde{S}_{n}(f)=\langle f_{n}^{\star},f\rangle_{\mathcal{H}^{d}}$. Then we obtain
@@
where Eqn. (79) is due to direct calculation on the derivative and in Eqn. (80) we use U-statistics law of large numbers [van2000asymptotic, Theorem 12.3]). Combining Eqn.(76), (78) and (81) directly yields that
@@ H.3 Proof of Lemma 5
In [fernandez2024general, Theorem 1 ], the similar result is shown under standard weak convergence instead of stable convergence. We aim to translate all the conditions and result into stable convergence setting.
@@
Using Lemma 1 again, we can see Condition (C1.1) is equivalent to the weak convergence of the sequence $\widetilde{S}_{n}(f)$ under any $F$ with $\mathbb{P}(F)>0$. Condition (C1.2) is not a probabilistic statement, so no modification is needed here. Also, Condition (C1.3) is modified according to $F$. With all these modifications, we can directly apply [fernandez2024general, Theorem 1 ] to get the stable convergence result. ∎
@@ H.5 Proof of Lemma 11
The difference between the statement of Lemma 11 and Proposition 6 in [fernandez2024general] is we only require convergence in probability in Eqn. (49) instead of almost sure convergence. However, it can be directly checked that the whole proof of Proposition 6 in [fernandez2024general] still goes through with this weaker condition.
@@
In the original proof of Proposition 6 in [fernandez2024general], the Statement (83) is derived used under almost sure convergence assumption. Therefore, the original proof is still valid under convergence in probability.
In order to apply Proposition 6 in [fernandez2024general] to prove Lemma 11, we pick $\mathcal{G}=\mathcal{H}$, $Q=\mathrm{id}$ and $L=K$ in the original statement. The whole proof goes through, which straightly implies Condition (C1.3) holds. ∎
@@ H.6 Proof of Lemma 12
We apply Le Cam’s first lemma [van2000asymptotic, Lemma 6.4] to prove this result. By Assumption 2, we have
@@
To further derive the asymptotic distribution of the estimator $\hat{\theta}_{n}$, we need to apply the Le Cam’s third lemma [van2000asymptotic, Theorem 6.6]. For simplicity, we denote
@@
where we use the Taylor expansion $\log(1+x)=x-x^{2}/2+O(x^{3})$ in the second line. To address remaining term, we first notice that by strong law of large numbers [durrett2019probability, Theorem 2.4.1], we have
@@
where $M_{2}^{h}=\mathrm{Var}_{\theta_{0}}[h(X)]$ and $\tau=\mathrm{Cov}_{\theta_{0}}[I(X,\theta_{0}),h(X)]$. Eqn. (86) is applied in the first line, and the standard CLT [durrett2019probability, Theorem 3.4.1] is used in the second line. By Le Cam’s third lemma, when under measure $\mathcal{L}_{n}$,
@@ Appendix I Auxiliary lemmas
By the central limit theorem of U-statistic [van2000asymptotic, Theorem 12.3], we have
@@
To be consistent with the notation in [niu2022reconciling], we use the notation $W_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$ to indicate that $W_{n}$ converges in distribution to a random variable $W$ conditionally on $\mathcal{F}_{n}$, i.e., for all $t\in\mathbb{R}$, at which the CDF of $W$ is continuous, we have
@@
Notice that $\mathbb{P}[W_{n}\leq t|\mathcal{F}_{n}]\leq 1$. We apply the bounded convergence theorem [durrett2019probability, Theorem 1.5.3] to obtain
@@
Lemma 21 (Marcinkiewicz-Zygmund strong law of large numbers, Theorem 2.5.12 and Theorem 2.5.13, [durrett2019probability]).
@@
We provide additional details on the order detection for the kernel exponential family in Section 5.2. The and null models and data generating processes in the two experiments are set as follows:
@@
    In the first setting, we set $p=2$ and $\theta_{1}=10$, $\theta_{2}$ is varied from 4-4 to $4$. The null model is the kernel exponential family with $p=1$ and $\theta_{1}\in\mathbb{R}$. Thus, when $\theta_{2}=0$, the data generating distribution is from the null model, otherwise it is from the alternative model.
@@
    In the second setting, we set $p=5$ to generate data. The parameters are varied for different instances and $\theta_{2},\ldots,\theta_{5}$ are randomly generated from $\mathrm{Unif}[-10,10]$. We aim to evaluate the power of the SKSD test when the size of null model increases.
@@
In this section, we provide additional details on the conditional Gaussian graphical model experiment in Section 5.3. The data generating process (DGP) is set as follows: for different $d$, we set $\gamma^{(1)}_{l}=2$ and $\gamma^{(2)}_{k}=-0.5$. For the interaction terms, we consider the 1D grid structure with reach $r=2$:
@@
where $\mathrm{vec}(\cdot)$ is the vectorization operator that stacks the columns of a matrix into a vector, and $b(x)=\sum_{k=1}^{d}\gamma_{k}^{(2)}x_{k}^{2}+\sum_{l=1}^{d}\gamma_{l}^{(1)}x_{i}\in\mathbb{R}$. Then, the estimators are given by
@@ Results.
In Figure 4, we present the power curves of SKSD tests under both multiplicative and additive local alternatives with different estimation procedures. When the observed data is generated from the null model, all tests control the Type-I error at the nominal level $0.05$. In the left panel of Figure 4, we consider the multiplicative local alternatives with data generated from generalized $\chi^{2}$ distributions. When $\alpha$ increases from $1$, the observed data become more heavy tailed, thus more distinguishable from the null model, and the power of both tests increases to nearly 1, which is consistent with the theoretical justification. In the right panel of Figure 4, we consider the additive local alternatives with data generated from the mixture of Gaussian distributions. When $w\in(0,1/2)$ increases, the observed data become more apparently multi-modal, thus the power of both tests increases, which again validates our theoretical results. In addition, we can see that the SKSD test with MLE has a slightly larger power than the one with minimum-KSD estimator in the multiplicative case, which confirms the influence of the estimation procedure on the power of the tests.
@@
Next, we consider the case when observed data is generalized $\chi^{2}$ distribution with power parameter $\alpha$. Similarly to the previous case, we expect the power of the tests to increase as $\alpha$ goes larger. Although the Type-I error is controlled at level $0.05$ for all four matrices, the MMD test does not show a consistent power curve as $\alpha$ increases. The other three tests with KSD, $W_{1}$, and K-S distance, show a similar consistent power curve, while the $W_{1}$ distance test has the largest power, followed by K-S distance and SKSD tests.

### v2

@@ Appendix E Preliminaries on stable convergence
Classical weak convergence results are established for the convergence of the quantile function, which is not sufficient for our purpose in this conditional convergence scenario, and a stronger notion of weak convergence is needed. To this end, we apply the notion of stable convergence [49], which is a stronger form of weak convergence and is suitable for studying the asymptotic behavior of conditional distributions.
@@
Generally speaking, stable convergence requires the weak limit of the sequence $Y_{n}$ keeps “stable” when conditioning on different sub-$\sigma$-algebra $\mathcal{G}$. The following lemma, adapted from Theorem 3.2 in [49], provides an equivalent characterization of stable convergence from this perspective.
@@
A key technical result applied in the proofs is the stable version of the central limit theorem. The following lemma, adapted from Theorem 6.1 in [49], states the conditions under which a martingale difference array converges stably to a normal distribution:
@@ F.1 Proof of Theorem 1
Under the regularity conditions, the law of large numbers for U-statistics [91, Theorem 12.3] implies that
@@ F.2 Proof of Theorem 2
Theorem 2 can be proved by following the strategy mentioned in the main text. We refer the audience to the proofs of Theorem 3 and Corollary 2 as natural extensions of this proof strategy.
@@ F.3 Proof of Theorem 3
Our proof relies on the concept of regular conditional distribution and convergence of Markov kernel, especially stable convergence. We provide introduction and basic results on the concept of stable convergence in Appendix E. See [49] for a detailed introduction to this framework.
We prove the desired stable convergence result through the following lemma, which is a generalization of [33, Theorem 1 ]. We first introduce some notation. Define $\mathcal{F}=\sigma(\cup_{n=1}^{\infty}\mathcal{F}_{n})$ to be the smallest $\sigma$-algebra generated by the whole data sequence. Notice that $\widetilde{T}_{n}$ and $\widetilde{S}_{n}$ both depend on the data through $\hat{\theta}_{n}$, which only depends $\{X_{i}\}_{i=1}^{n}$. Based on this observation, deriving the convergence results conditioning on $\mathcal{F}$ is equivalent to conditioning on $\mathcal{F}_{n}$. Thus, for simplicity, we work on the $\mathcal{F}$-stable convergence of the test statistic in the proof in this part.
@@
We use the definition of stable convergence in [49, Page 6 ] to prove it. Recall Definition of the operator norm in Eqn. (30). Let us choose
@@ F.4 Proof of Theorem 4
We use the results in Theorem 2 and 3 combining with [71, Lemma 3 ] to prove the asymptotic validity of the test. We state their lemma here for completeness.
@@
which is equivalent to $n\widetilde{T}_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$ with notation used in [71]. Applying Lemma 19, we have
@@
where in the first inequality, we use the fact that $\mathrm{KSD}(\mathcal{L}_{\hat{\theta}_{n}},\mathcal{L})\geq\inf_{\theta\in\Theta}\mathrm{KSD}(\mathcal{L}_{\theta},\mathcal{L})=2\varepsilon$. The last expression can be bounded by the following lemma, which is indeed a stronger result. The proof of this result can be found in [66, Lemma 2 ].
@@ F.5 Proof of Proposition 2
We will verify the conditions in the Kolmogorov’s three series theorem [27, Page 150]. By choosing the truncation threshold $A$ in the theorem to be $\infty$, the first condition in the three series holds automatically. Thus, it suffices to show that
@@ F.6 Proof of Corollary 2
We directly prove Eqn. (37) in this part. [33, Theorem 1 ] is used to prove this result. We state it here for convenience.
@@
still hold for any $\varepsilon>0$. Then by standard Lindeberg-Feller Central Limit Theorem [27, Theorem 3.4.10 ], we verify Condition (C2.1).
@@ F.7 Verification of the conditions in Lemma 5
We will mainly use a slightly modified version of [33, Proposition 6 ]. The version of usage is stated as below:
@@ G.2 Verification of conditions in Lemma 14
By Le Cam’s third lemma [91, Theorem 6.6], we conclude that under local alternative $\mathcal{L}_{n}$,
@@ G.3 Proof of the additive case in Theorem 5
$\mu_{i}$ $=h\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})],$
@@
$p_{1,n}(x)$ $=\left(1-\frac{h}{\sqrt{n}}\right)p_{\theta_{0}}(x)+\frac{h}{\sqrt{n}}g(x)$
$=p_{\theta_{0}}(x)\left(1+\frac{h}{\sqrt{n}}\left(\frac{g(x)}{p_{\theta_{0}}(x)}-1\right)\right),$
which can be fitted into the multiplicative local alternatives with $h(x)=h(g(x)/p_{\theta_{0}}(x)-1)$. It can be directly checked that Assumption 2 hold. Thus, we immediately have Theorem 5 holds with
$\mu_{i}$ $=\mathbb{E}_{\theta_{0}}\left[S^{*}(\phi_{i},X,\theta_{0})\cdot h\left(\frac{g(x)}{p_{\theta_{0}}(x)}-1\right)\right]$
$=h\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})]-h\mathbb{E}_{\theta_{0}}[S^{*}(\phi_{i},X,\theta_{0})]$
$=h\mathbb{E}_{\mathcal{L}_{g}}[S^{*}(\phi_{i},X,\theta_{0})],$
@@ H.1 Proof of Lemma 3
By noticing that $X\sim\mathbb{P}_{\theta}$ and the property of kernelized Stein class in Eqn. (6), we have
@@ H.2 Proof of Lemma 4
We begin by explicitly calculate the left hand side of Equation (73). It is straightforward to check that $\widetilde{S}_{n}(f)$ is a linear functional on $\mathcal{H}^{d}$ By Riesz representation theorem [101, Page 81 ], there exists a unique $f_{n}^{\star}\in\mathcal{H}^{d}$ such that $\widetilde{S}_{n}(f)=\langle f_{n}^{\star},f\rangle_{\mathcal{H}^{d}}$. Then we obtain
@@
where Eqn. (79) is due to direct calculation on the derivative and in Eqn. (80) we use U-statistics law of large numbers [91, Theorem 12.3]). Combining Eqn.(76), (78) and (81) directly yields that
@@ H.3 Proof of Lemma 5
In [33, Theorem 1 ], the similar result is shown under standard weak convergence instead of stable convergence. We aim to translate all the conditions and result into stable convergence setting.
@@
Using Lemma 1 again, we can see Condition (C1.1) is equivalent to the weak convergence of the sequence $\widetilde{S}_{n}(f)$ under any $F$ with $\mathbb{P}(F)>0$. Condition (C1.2) is not a probabilistic statement, so no modification is needed here. Also, Condition (C1.3) is modified according to $F$. With all these modifications, we can directly apply [33, Theorem 1 ] to get the stable convergence result. ∎
@@ H.5 Proof of Lemma 11
The difference between the statement of Lemma 11 and Proposition 6 in [33] is we only require convergence in probability in Eqn. (49) instead of almost sure convergence. However, it can be directly checked that the whole proof of Proposition 6 in [33] still goes through with this weaker condition.
@@
In the original proof of Proposition 6 in [33], the Statement (83) is derived used under almost sure convergence assumption. Therefore, the original proof is still valid under convergence in probability.
In order to apply Proposition 6 in [33] to prove Lemma 11, we pick $\mathcal{G}=\mathcal{H}$, $Q=\mathrm{id}$ and $L=K$ in the original statement. The whole proof goes through, which straightly implies Condition (C1.3) holds. ∎
@@ H.6 Proof of Lemma 12
We apply Le Cam’s first lemma [91, Lemma 6.4] to prove this result. By Assumption 2, we have
@@
To further derive the asymptotic distribution of the estimator $\hat{\theta}_{n}$, we need to apply the Le Cam’s third lemma [91, Theorem 6.6]. For simplicity, we denote
@@
where we use the Taylor expansion $\log(1+x)=x-x^{2}/2+O(x^{3})$ in the second line. To address remaining term, we first notice that by strong law of large numbers [27, Theorem 2.4.1], we have
@@
where $M_{2}^{h}=\mathrm{Var}_{\theta_{0}}[h(X)]$ and $\tau=\mathrm{Cov}_{\theta_{0}}[I(X,\theta_{0}),h(X)]$. Eqn. (86) is applied in the first line, and the standard CLT [27, Theorem 3.4.1] is used in the second line. By Le Cam’s third lemma, when under measure $\mathcal{L}_{n}$,
@@ Appendix I Auxiliary lemmas
By the central limit theorem of U-statistic [91, Theorem 12.3], we have
@@
To be consistent with the notation in [71], we use the notation $W_{n}|\mathcal{F}_{n}\overset{d,p}{\to}W$ to indicate that $W_{n}$ converges in distribution to a random variable $W$ conditionally on $\mathcal{F}_{n}$, i.e., for all $t\in\mathbb{R}$, at which the CDF of $W$ is continuous, we have
@@
Notice that $\mathbb{P}[W_{n}\leq t|\mathcal{F}_{n}]\leq 1$. We apply the bounded convergence theorem [27, Theorem 1.5.3] to obtain
@@
Lemma 21 (Marcinkiewicz-Zygmund strong law of large numbers, Theorem 2.5.12 and Theorem 2.5.13, [27]).
@@
We provide additional details on the order detection for the kernel exponential family in Section 5.2. The null models and data generating processes in the two experiments are set as follows:
@@
    In the first setting, we set $r=2$ and $\theta_{1}=10$, $\theta_{2}$ is varied from 4-4 to $4$. The null model is the kernel exponential family with $r=1$ and $\theta_{1}\in\mathbb{R}$. Thus, when $\theta_{2}=0$, the data generating distribution is from the null model, otherwise it is from the alternative model.
@@
    In the second setting, we set $r=5$ to generate data. The parameters are varied for different instances and $\theta_{2},\ldots,\theta_{5}$ are randomly generated from $\mathrm{Unif}[-10,10]$. We aim to evaluate the power of the SKSD test when the size of null model increases.
@@
In this section, we provide additional details on the conditional Gaussian graphical model experiment in Section 5.3. The data generating process is set as follows: for different $d$, we set $\gamma^{(1)}_{l}=2$ and $\gamma^{(2)}_{k}=-0.5$. For the interaction terms, we consider the 1D grid structure with reach $r=2$:
@@
where $\mathrm{vec}(\cdot)$ is the vectorization operator that stacks the columns of a matrix into a vector, and $b(x)=\sum_{k=1}^{d}\gamma_{k}^{(2)}(x^{(k)})^{2}+\sum_{l=1}^{d}\gamma_{l}^{(1)}x^{(l)}\in\mathbb{R}$. Then, the estimators are given by
@@ Results.
In Figure 4, we present the power curves of SKSD tests under both multiplicative and additive local alternatives with different estimation procedures. When the observed data are generated from the null model, all tests control the Type-I error at the nominal level $0.05$. In the left panel of Figure 4, we consider the multiplicative local alternatives with data generated from generalized $\chi^{2}$ distributions. When $\alpha$ increases from $1$, the observed data become more heavy tailed, thus more distinguishable from the null model, and the power of both tests increases to nearly 1, which is consistent with the theoretical justification. In the right panel of Figure 4, we consider the additive local alternatives with data generated from the mixture of Gaussian distributions. When $w\in(0,1/2)$ increases, the observed data become more apparently multi-modal, thus the power of both tests increases, which again validates our theoretical results. In addition, we can see that the SKSD test with MLE has a slightly larger power than the one with minimum-KSD estimator in the multiplicative case, which confirms the influence of the estimation procedure on the power of the tests.
@@
Next, we consider the case when observed data are generalized $\chi^{2}$ distribution with power parameter $\alpha$. Similarly to the previous case, we expect the power of the tests to increase as $\alpha$ goes larger. Although the Type-I error is controlled at level $0.05$ for all four matrices, the MMD test does not show a consistent power curve as $\alpha$ increases. The other three tests with KSD, $W_{1}$, and K-S distance, show a similar consistent power curve, while the $W_{1}$ distance test has the largest power, followed by K-S distance and SKSD tests.